a group of 7 members having a majority of boys is to be formed out of 6 boys and 4 girls. the number of ways the group can be formed

Answer by stanbon(75887)

a group of 7 members having a majority of boys is to be formed out of 6 boys and 4 girls. the number of ways the group can be
------
6boys/1 girl: 6C6*4C1 = 1*4 = 4
5boys/2 girls: 6C5*4C2 = 6*6 = 36
4boys/3 girls: 6C4*4C3 = 15*4 = 60
----
Ans: Total = 100 ways
==================
Cheers,
Stan H.
==================